body of mass 200 g executing simple
harmonic motion has a velocity of 3 cms^-1
when its displacement is 4 cm and a velocity
of 4 cms^-1when the displacement is 3 cm.
Calculate
(i) amplitude of oscillation, and
(ii) period of oscillation.
1
Expert's answer
2019-05-24T09:57:57-0400

i) Velocity of the particle:


"v^2=\omega^2(A^2-x^2)"

So


"4^2=\omega^2(A^2-3^2)"

"3^2=\omega^2(A^2-4^2)"

"\frac{(A^2-3^2)}{(A^2-4^2)}=\frac{16}{9}"

"A=5\ cm"

ii)


"4^2=\omega^2(5^2-3^2)"

"\omega=1\frac{rad}{s}"

"T=\frac{2\pi}{\omega}=\frac{2\pi}{1}=6.28\ s"